how many grams of lactose must be added to 655 grams of water in order to prepare a solution with a percent mass concentration of 10%?


1
Expert's answer
2022-02-11T12:29:02-0500

Let m "=" The mass of lactose


"\frac{m}{655+m}=\frac{10}{100}"


"100m=6550+10m"

"90m=6550"

"m=72.78g"